Suppose that the population in question is the set of all registered voters in the Philippines....

  1. Suppose that the population in question is the set of all registered voters in the Philippines. A rough estimate of the total number of registered voters is 90,000,000. Calculate the random samples where the error of tolerance is 0.01%

  1. In a survey on the level of awareness of registered voters registered voters regarding the computerization of the election processes resulted into:

(5)

(4)

(3)

(2)

(1)

Strongly Agree

Agree

Undecided

Disagree

Strongly Disagree

Level of Awareness

800

200

50

100

50

compute the mean and interpret the results.
Answer #1

Total level of awareness = 800+200+50+100+50 = 1200

Arranging the data we get

strongly agree agree undecided disagree strongly disagree
x 5 4 3 2 1

P(x)

Level of

Awareness

800/1200 200/1200 50/1200 100/1200 50/1200

Mean = sum(Pi*Xi)

= 800/1200 * 5 + 200/1200 * 4 + 50/1200*3 + 100/1200 *2 + 50/1200* 1

= 800/1200 * 5 + 200/1200 * 4 + 50/1200*3 + 100/1200 *2 + 50/1200* 1

= 4.33

Thus, mean indicate that average level of awareness is between agree and strongly agree, and is inclined towards agree